# Energy consumption of code small language models serving with runtime engines and execution providers

### v1

@@ 4.1 Research goal and questions
The research goal [49] of this study is to analyze deep learning runtime engines and execution providers for the purpose of understanding their impact with respect to energy consumption, execution time, and computing-resource utilization from the point of view of software engineers conducting inference in the context of code SLMs.
@@
    How is the energy consumption affected by the execution provider and runtime engine used for inference from code SLMs?
@@
    How is the execution time affected by the execution provider and runtime engine used for inference from code SLMs?
@@
    How is the computing-resource utilization (CPU usage, RAM usage, GPU usage, GPU memory utilization, and GPU used memory) affected by the execution provider and runtime engine used for inference from code SLMs?
@@ 4.2 Experimental units
In this subsection, we detail the experimental units and the process used to construct the input dataset for our study. The experimental units refer to the selected code SLMs used for inference, while the dataset construction draws from the widely-used HumanEval dataset [23]. The experimental units were subjected to treatments in the form of different deep learning serving configurations to evaluate them during inference.
The selection process for the code generation SLMs in this study was conducted with a systematic and structured approach [50], reflecting the state of available models as of November 6th, 2024. Given the time required for executing experiments, and analyzing results, this study might not include other recent code SLMs after such date.
Using the Hugging Face API, metadata for 104,816 text-generation decoder models was retrieved, including detailed model card information. A series of filters were applied to refine the selection: models exceeding 5 billion parameters were excluded to focus on SLMs, and priority was given to models with potential HumanEval results explicitly mentioned in their metadata, model cards, or flagged as gated models. Only base models, identified within their respective model trees, were retained, narrowing the pool to 340 models. Manual verification ensured that the final 29 models were genuinely base models, free from mislabeled fine-tuned or quantized versions, and had HumanEval evaluation results to confirm their suitability for code-related tasks. From this pool, 12 code SLMs, presented in Table 1, were randomly selected to ensure diversity and representativeness. A replication package containing the extracted data is available for reproducibility and future research (see data availability statement in Section 1).
Table 1: Overview of the 12 Code SLMs used for inference in this study. These models were selected through a systematic and reproducible process (see also 4.2).
Model Size Category Model Release Date Size [M/B] Affiliation HumanEval (Pass@1) Hugging Face Name
@@ 4.3 Dataset construction
Our input dataset was derived utilizing a methodology similar to those outlined by Fried et al. and Bavarian et al. [51, 52]. Specifically, we selected a subset of prompts from the HumanEval dataset [23] by taking each problem and processing it as follows: we tokenized the entire problem and then randomly selected a sequence of 10 to 15 tokens. This token sequence was decoded and used as an input prompt for the model, with the aim of generating additional tokens based on the sampled input. This method enabled us to create a dataset that was both controlled to fit our computational setup and diverse enough to effectively evaluate the small language models (SLMs) in code generation tasks with respect to our independent variables.
@@ 4.4 Study design
We conduct a technology oriented experiment [53] following the guidelines by Jedlitschka et al. [54]. We selected twelve trained models and structured our investigation into a multi-stage pipeline, as illustrated in Figure 2. This pipeline consists of five main stages [55, 56]:
@@
    Data management: This foundational phase is critical in setting up our experiment to test code generation models. We start with the HumanEval dataset [23], which includes 164 coding problems, and adapt it to better suit the specific requirements of our study involving SLMs. This adaptation involves generating a new dataset from the HumanEval problems for the code generation task (see also 4.3).
@@
    Modelling: Our study focuses on analyzing different configurations of deep learning serving infrastructures (see subsection 4.5.1), utilizing pre-trained models. This approach is in alignment with the objectives of our study (see also 4.2).
@@ 4.5.1 Independent variables
In this study we define one independent variable, Serving Configuration, which is a duplet of a runtime engine and an execution provider (see section 2). The setup consists of evaluating different deep learning serving configurations. We initially obtain eight possible configurations by combining two execution providers (CPU and CUDA) with four runtime engines (TORCH, ONNX, OV, and JIT). However, we exclude the $<$OV, CUDA$>$ configuration, as OV is specifically optimized for Intel hardware and does not have compatibility with the CUDA execution provider [57]. As a result, we analyze seven configurations, each representing a unique combination of runtime engine and execution provider (e.g., $<$ONNX, CUDA$>$).
@@ 4.5.2 Dependent variables
To measure the impact of our independent variable to energy efficiency, and performance efficiency, the study tracks some of the most widely measured and relevant metrics in the field according to the Green Software Measurement Model [58]: the energy consumption, execution time, and computing-resource utilization (CPU usage, RAM usage, GPU usage, GPU memory utilization, and GPU used memory) during the inference (see also Table 2).
@@
{tblr}
row1 = c, cell21 = c, cell31 = r=7c, hline1-3,10 = -, hline4-9 = 2-5, Class & Name Description Scale Operationalization
@@
provider: $<$TORCH, CPU$>$, $<$ONNX, CPU$>$,
$<$OV, CPU$>$, $<$JIT, CPU$>$, $<$TORCH, CUDA$>$,
$<$ONNX, CUDA$>$, and $<$JIT, CUDA$>$. Nominal See sections 2 and 4.5.
@@
Execution time Average of duration to do inference with test
@@ 4.6 Experiment setting
All these experiments have been conducted with an NVIDIA GeForce RTX 4090 GPU with 24 GB memory with CUDA Version 12.4, and an AMD Ryzen 9 7950X 16-Core Processor CPU, 2.5 GHz frequency.
The complete experiment and data analysis is implemented in Python 3.10.12, using HuggingFace Transformers 4.40.2 [59], and Optimum 1.19.2666https://huggingface.co/docs/optimum/. To implement the serving configuration, we followed the implementations and model exportation provided by HuggingFace.
To accurately measure the variables outlined in our study, we use the energibridge profiler, nvidia-smi profiler, and a wattmeter. These tools are recognized in the scientific community for its effectiveness in assessing energy consumption of computational processes, as evidenced by its application in previous research [60, 61]. EnergiBridge was chosen for its simplicity and ready-to-use setup, enabling efficient data collection across various platforms without extensive configuration. Nvidia-smi, optimized for NVIDIA GPUs, was the ideal tool for measuring GPU-specific energy consumption and utilization. Finally, the wattmeter provided direct and precise measurements of total device energy consumption. It was used to verify that the energy reported by the CPU and GPU profilers matched the total energy consumption measured by the wattmeter, ensuring the accuracy and validity of our energy measurements.
Energibridge. [^7] https://github.com/tdurieux/EnergiBridge EnergiBridge is a cross-platform tool designed to measure software energy consumption on various operating systems (Linux, Windows, and macOS) and CPU architectures (Intel, AMD, and Apple ARM). Developed in Rust, it simplifies energy monitoring by providing a unified interface and collects data through low-level system calls, converting measurements into watts for standardized reporting across different systems [62].
@@
Wattmeter. [^9] https://vitriko.eu/regleta-inteligente-netio-powerbox-4kf The wattmeter is a tool for measuring electrical power. In our experimental setup we used the NETIO PowerBOX 4KF, an intelligent power strip with four individually controllable outlets. This device is integrated into our system to monitor and record key electrical metrics such as energy consumption. We retrieve data through its JSON-based API.
@@ 4.7 Experiment execution
In this study, we deploy a server that receives HTTP requests from a client. The server hosts the developed API, which processes input prompts and returns the model’s output, specifically the generated decoded tokens. The experiment involves selecting one of the seven serving configurations and performing the following steps for each configuration: for each model, the model is loaded, and inference is conducted on all 164 requests using the input dataset. During the inference process for each serving configuration, we collect the dependent variables using the three profilers described in section 4.6. To ensure stability, we run the server for five minutes before starting the inferences and for another five minutes after completing the inferences with each serving configuration. The experiment is repeated 10 times to ensure consistency in the results, as this number of repetitions was sufficient to observe minimal variability and provide reliable medians across configurations. Still, each repetition of the experiment involves 7 serving configurations, 12 models, and 164 input requests per model, resulting in a substantial amount of data collected.
@@ 4.8 Data analysis
{tblr}
cell23 = r, cell33 = r, cell41 = r=5c, cell43 = r, cell53 = r, cell63 = r, cell73 = r, cell83 = r, hline1-4,9 = -, hline5-8 = 2-4, RQs & Variable Statistical test/
Post-hoc test
Null hypothesis
@@ 5 Results
In this section, we report the results of our experiment, addressing our research questions and highlighting key takeaways. We initially evaluated seven deep learning serving configurations. Following the experiments, we reviewed the outputs of the scripts and inferences, identifying error-prone responses in the TorchScript configurations. As a result, these configurations were excluded from further analysis. The results presented here focus on comparisons across the final five selected configurations while using results from twelve different code SLMs, as shown in Tables 4 - 5 and Figures 3 - 5 (one color for each model in the box plots).
Due to the number of used models for each serving configuration we report results aggregated by model size category (see table 1) for all metrics. The categories are organized based on model size category, with category 1 representing the smallest models and category 3 representing the largest models. We present more detailed results for all evaluation metrics in the replication package.
5.1 Energy consumption of code SLMs for code generation (RQ1).
@@
The median energy consumption for configurations using the CUDA execution provider is consistently lower than that of the CPU execution provider. Notably, the $<$TORCH, CUDA$>$ configuration stands out as the most energy-efficient across all models, a trend that is also evident in the aggregated results.
The $<$ONNX, CUDA$>$ configuration exhibits the highest median energy consumption values among the CUDA execution provider configurations. Similarly, $<$TORCH, CPU$>$ configurations consume the most energy across the majority of models when using the CPU execution provider. This pattern is also reflected in the aggregated results.
ONNX optimizations are significantly more effective when paired with the CPU execution provider, demonstrating better energy efficiency compared to the deep learning framework alone. Among the optimized runtime engines, ONNX, along with OV, consumes less energy, with ONNX emerging as the most energy-efficient in CPU execution provider configurations. However, when using the CUDA execution provider, ONNX fails to provide the same level of energy savings, as it does not outperform the baseline TORCH runtime engine, which remains the most energy-efficient. This discrepancy in energy performance likely stems from the way the ONNX runtime engine interacts with each execution provider.
While using the CPU execution provider, it was observed that the extended execution times (see Table 4) led to a significant increase in energy consumption by the CPU. This prolonged inference time, especially with more complex models, may cause CPU configurations to use substantially more energy. The inefficiency of the CPU execution provider in energy conservation highlights the need for optimized resource allocation and management in deep learning serving environments to minimize unnecessary energy consumption. In contrast, CUDA configurations, despite utilizing both CPU and GPU resources, were far more energy-efficient for intensive tasks due to their significantly faster execution times.
@@
Answer to RQ1: Energy consumption is significantly affected by the serving configuration, with CUDA configurations being more efficient than CPU configurations. $<$TORCH, CUDA$>$ provides the best balance between energy efficiency and reliability, while optimized runtime engines like OV and ONNX, particularly ONNX, reduce energy consumption in CPU configurations but are less effective with CUDA compared to TORCH.
@@
1 $<$ONNX, CPU$>$ 45069.35 655.09
1 $<$OV, CPU$>$ 58374.88 561.61
1 $<$TORCH, CPU$>$ 64993.75 773.27
2 $<$ONNX, CPU$>$ 178641.05 2100.83
2 $<$OV, CPU$>$ 223821.62 1355.18
2 $<$TORCH, CPU$>$ 216854.71 2583.76
3 $<$ONNX, CPU$>$ 393187.23 3958.18
3 $<$OV, CPU$>$ 400527.50 2300.14
3 $<$TORCH, CPU$>$ 512731.24 5668.91
1,2,3 $<$ONNX, CPU$>$ 205632.54 2238.03
1,2,3 $<$OV, CPU$>$ 227574.67 1405.64
1,2,3 $<$TORCH, CPU$>$ 264859.90 3008.65
@@
1 $<$ONNX, CUDA$>$ 13404.99 20692.81 34138.28 277.12
1 $<$TORCH, CUDA$>$ 5193.32 11250.87 16432.53 124.03
2 $<$ONNX, CUDA$>$ 32113.95 55652.21 87788.33 630.64
2 $<$TORCH, CUDA$>$ 7287.23 24247.47 31525.72 168.74
3 $<$ONNX, CUDA$>$ 39012.21 81516.82 120509.17 763.59
3 $<$TORCH, CUDA$>$ 12196.07 50052.11 62301.62 260.62
1,2,3 $<$ONNX, CUDA$>$ 28177.05 52620.61 80811.93 557.12
1,2,3 $<$TORCH, CUDA$>$ 8225.54 28516.82 36753.29 184.46
5.2 Execution time of code SLMs for code generation (RQ2).
@@
Extreme cases are observed with models such as phi2 using $<$TORCH, CPU$>$, where the median execution time (5983.52 seconds) is more than 20 times longer than with the $<$TORCH, CUDA$>$ configuration (251.97 seconds), highlighting significant bottlenecks in CPU-based processing for complex computational tasks. Similar extreme cases are noted within the largest model size category (Category 3) using TORCH, where the CPU execution provider’s execution time (5668.91 seconds) exceeds the CUDA execution provider’s time (260.62 seconds) by more than 20 times.
@@
Welch’s ANOVA confirmed significant differences in execution times across different runtime engines. The subsequent Games-Howell post-hoc tests annotate that all configurations are statistically distinct from one another, confirming the marked impact of the execution provider and runtime engine choice on the efficiency of computational tasks (H.2.0). The only exceptions were that the execution times of the $<$ONNX, CPU$>$ and $<$TORCH, CPU$>$ configurations did not differ significantly for three out of the twelve models (bloomz-560m, codeparrot-small, and tiny-starcoder), resulting in the rejection of H.2.0 for only these models.
Answer to RQ2: Execution time is significantly influenced by the serving configuration, with CUDA configurations consistently outperforming CPU configurations. $<$TORCH, CUDA$>$ delivers the best performance, while OV and ONNX optimize execution time, especially OV, with CPU execution provider but do not achieve comparable improvements with CUDA execution provider, reflecting similar patterns seen in energy consumption.
5.3 Computing-resource utilization of code SLMs for code generation (RQ3).
@@
For statistical evaluation, different tests were applied depending on the normality and variance homogeneity of the data. Welch’s ANOVA was used for data with unequal variances but normal distributions, while the Kruskal-Wallis test was employed for non-normally distributed data. The significance of differences between configurations was assessed using these tests, with p-values indicating significant differences with values below $\alpha=0.05$. Furthermore, post-hoc analyses using Games-Howell or Dunn’s test were performed to identify specific pairwise differences. In terms of computing-resource utilization metrics, all serving configurations showed significant differences from one another (H.3.1.0, H.3.2.0, H.3.3.0, H.3.4.0, and H.3.5.0), with a few exceptions. Specifically, RAM usage between certain serving configuration pairs—such as $<$OV, CPU$>$ and $<$TORCH, CPU$>$, $<$ONNX, CUDA$>$ and $<$TORCH, CUDA$>$, $<$OV, CPU$>$ and $<$ONNX, CPU$>$, as well as $<$TORCH, CPU$>$ and $<$ONNX, CUDA$>$—did not exhibit significant differences in two out of the twelve models for each pair. As a result, H.3.2.0 was rejected for those specific cases.

### v2

@@ 4.1 Research goal and questions
The research goal [64] of this study is to analyze deep learning runtime engines and execution providers for the purpose of understanding their impact with respect to resource utilization in terms of energy consumption, execution time, and computing-resource utilization from the point of view of software engineers conducting inference in the context of code generation SLMs.
@@
    How is the energy consumption affected by the execution provider and runtime engine used for inference from code generation SLMs?
@@
    How is the execution time affected by the execution provider and runtime engine used for inference from code generation SLMs?
@@
    How is the computing-resource utilization (CPU usage, RAM usage, GPU usage, GPU memory utilization, and GPU used memory) affected by the execution provider and runtime engine used for inference from code generation SLMs?
@@ 4.2 Experimental units
In this subsection, we detail the experimental units and the process used to construct the input dataset for our study. The experimental units refer to the selected code generation SLMs used for inference, while the dataset construction draws from the widely-used HumanEval dataset [25]. The experimental units were subjected to treatments in the form of different deep learning serving configurations to evaluate them during inference.
The selection process for the code generation SLMs in this study was conducted with a systematic and structured approach [65], reflecting the state of available models as of November 6th, 2024. Given the time required for executing experiments, and analyzing results, this study might not include other recent code generation SLMs after such date.
Using the Hugging Face API, metadata for 104,816 text-generation decoder models was retrieved, including detailed model card information. A series of filters were applied to refine the selection: models exceeding 5 billion parameters were excluded to focus on SLMs, and priority was given to models with potential HumanEval results explicitly mentioned in their metadata, model cards, or flagged as gated models. Only base models, identified within their respective model trees, were retained, narrowing the pool to 340 models potential candidates.
Afterwards, through metadata inspection and structured filtering based on the explicit mention and reporting of HumanEval benchmarks [25], we narrowed down the selection to 29 pretrained base models. ‘potential HumanEval results’ refers specifically to models whose metadata explicitly indicated having undergone HumanEval evaluation, whether in structured through the HuggingFace API or textual forms. We manually verified the model metadata to ensure that the selected models are base pretrained models, not subjected to quantization or fine-tuning.
Given practical constraints, such as total experimental runtime and available infrastructure resources, we estimated experiment durations and concluded that exhaustively testing all 29 models across all configurations was infeasible. Consequently, we randomly selected 12 models, presented in Table 1, intentionally ensuring diversity and representativeness of different sizes (e.g., 110 [M], 1 [B], and 3[B]), affiliation (e.g., academia and industry), and release dates (from 2021 to 2024). A replication package containing the extracted data is available for reproducibility and future research (see data availability statement in Section 1).
Our primary focus remained on understanding the impact of the serving infrastructure rather than extensively characterizing individual models. Nevertheless, we confirmed through preliminary tests that model outputs remained consistent across configurations, except for JIT configurations, whose inconsistencies were explicitly investigated and documented within our publicly available replication package.
Table 1: Overview of the 12 code generation SLMs used for inference in this study. These models were selected through a systematic and reproducible process (see also 4.2).
Model Size
Category
Model Release Date Size [M/B] Affiliation
HumanEval
(Pass@1)
Hugging Face
Name
@@ 4.3 Dataset construction
Our input dataset was derived utilizing a methodology similar to those outlined by Fried et al. and Bavarian et al. [66, 67]. Specifically, we selected a subset of prompts from the HumanEval dataset [25] by taking each problem and processing it as follows: we tokenized the entire problem and then randomly selected a sequence of 10 to 15 tokens to form an input prompt. This token sequence was decoded and used as an input prompt for the model, with the aim of generating additional tokens based on the sampled input. By repeating this procedure across the entire HumanEval set, we created a collection of short‐prompt (fewer than 15 tokens) completion requests that are both controlled to fit our computational setup and diverse enough to effectively evaluate the SLMs in code generation tasks with respect to our dependent variables [60, 61]. The resulting dataset closely emulates interactive IDE‐style auto‐complete code task and ensures uniformity in prompt length when comparing inference efficiency across different serving configurations [31].
Although our dataset is generated from the HumanEval benchmark, consistent with prior work that evaluates language models on this suite [4], we concentrate exclusively on auto-complete code scenario rather than large-context or conversational code generation scenarios. Modern IDEs typically trigger completion suggestions after only a few tokens have been entered, making a 10 to 15 token window a realistic proxy for real-world usage [68]. Consequently, our dataset captures short-prompt use cases and does not cover tasks that involve navigating a large codebase, synthesizing entire functions across multiple files, or engaging in multi-turn conversational exchanges with an LLM, such as understanding how functions in different files connect or spotting a small bug in code evolving over time [31, 68]. While this choice restricts our findings to auto-complete task rather than full-function or deep conversational tasks, it provides a controlled environment for rigorously comparing serving behavior, such as execution time, energy consumption, and computing-resource utilization, when generating small code completions. All data and scripts for dataset creation are available in our replication package.
@@ 4.4 Study design
We conduct a technology-oriented experiment [33] following the guidelines by Jedlitschka et al. [69]. We selected twelve trained models and structured our investigation into a multi-stage pipeline, as illustrated in Figure 2. This pipeline consists of five main stages [70, 71]:
@@
    Data management: This foundational phase is critical in setting up our experiment to test code generation SLMs. We start with the HumanEval dataset [25], which includes 164 coding problems, and adapt it to better suit the specific requirements of our study involving SLMs. This adaptation involves generating a new dataset through a token sampling strategy from HumanEval problems for the code generation task (see also 4.3). Our study design leverages these generated prompts to evaluate multiple serving configurations, emulating a practical yet controlled environment. While HumanEval provides self-contained problems that reflect basic to intermediate coding challenges, it does not encompass, for instance, large-scale or multi-file code generation. We selected it for consistency with existing literature and ease of reproducibility [4].
@@
    Modelling: Our study focuses on analyzing different configurations of deep learning serving infrastructures (see Subsection 4.5.1), utilizing pre-trained models. This approach is in alignment with the objectives of our study (see also Subsection 4.2).
@@ 4.5.1 Independent variables
In this study we define one independent variable, Serving Configuration, which is a duplet of a runtime engine and an execution provider (see Section 2). The setup consists of evaluating different deep learning serving configurations. We initially obtain eight possible configurations by combining two execution providers (CPU and CUDA) with four runtime engines (TORCH, ONNX, OV, and JIT). However, we exclude the $\langle OV,CUDA\rangle$ configuration, as OV is specifically optimized for Intel hardware and does not have compatibility with the CUDA execution provider [72]. As a result, we analyze seven configurations, each representing a unique combination of runtime engine and execution provider (e.g., $\langle ONNX,CUDA\rangle$).
@@ 4.5.2 Dependent variables
To measure the impact of our independent variable to resource utilization, the study tracks some of the most widely measured and relevant metrics in the field according to the Green Software Measurement Model [73]: the energy consumption, execution time, and computing-resource utilization (CPU usage, RAM usage, GPU usage, GPU memory utilization, and GPU used memory) during the inference (see also Table 2).
@@
Class Name Description Scale Operationalization
@@
provider: $\langle TORCH,CPU\rangle$, $\langle ONNX,CPU\rangle$,
$\langle OV,CPU\rangle$, $\langle JIT,CPU\rangle$, $\langle TORCH,CUDA\rangle$,
$\langle ONNX,CUDA\rangle$, and $\langle JIT,CUDA\rangle$. Nominal See sections 2 and 4.5.
@@
Execution time Average of duration to do inference with input
@@ 4.6 Experiment setting
All these experiments have been conducted with an NVIDIA GeForce RTX 4090 GPU with 24 GB memory with CUDA Version 12.4, and an AMD Ryzen 9 7950X 16-Core Processor CPU, 2.5 GHz frequency with 32 GB of DDR5 5600 MHz RAM.
The complete experiment and data analysis is implemented in Python 3.10.12, using HuggingFace Transformers 4.40.2 [74], and Optimum 1.19.2777https://huggingface.co/docs/optimum/. To implement the serving configuration, we followed the implementations and model exportation provided by HuggingFace.
To accurately measure the variables outlined in our study, we use the energibridge profiler, nvidia-smi profiler, and a wattmeter. These tools are recognized in the scientific community for their effectiveness in assessing the resource utilization of computational processes, as evidenced by their application in previous research [75, 76]. EnergiBridge was chosen for its simplicity and ready-to-use setup, enabling efficient data collection across various platforms without extensive configuration. Nvidia-smi, optimized for NVIDIA GPUs, was the ideal tool for measuring GPU-specific resource utilization. The wattmeter provided direct and precise measurements of total device energy consumption. It was used to verify that the energy reported by the CPU and GPU profilers matched the total energy consumption measured by the wattmeter, ensuring the accuracy and validity of our energy measurements. Finally, we used the Linux date command to record timestamps at the start and end of each experiment. These were cross-referenced with the timestamps recorded by the profiling tools to ensure alignment and to verify that the system was either idle or actively running inferences.
EnergiBridge. [^8] https://github.com/tdurieux/EnergiBridge EnergiBridge is a cross-platform tool designed to measure software resource utilization, focused on energy consumption, on various operating systems (Linux, Windows, and macOS) and CPU architectures (Intel, AMD, and Apple ARM). Developed in Rust, it simplifies resources monitoring by providing a unified interface and collects data through low-level system calls, converting measurements into watts for standardized reporting across different systems [77].
@@
Wattmeter. [^10] https://vitriko.eu/regleta-inteligente-netio-powerbox-4kf The wattmeter is a tool for measuring electrical power. In our experimental setup, we used the NETIO PowerBOX 4KF, an intelligent power strip with four individually controllable outlets. The experimental system was connected to the monitored outlets, enabling accurate measurement of key electrical metrics such as energy consumption. Measurements were retrieved through the device’s JSON-based API. The wattmeter was configured according to the manufacturer’s setup instructions to ensure consistency and reproducibility in all experiments.
@@ 4.7 Experiment execution
In this study, we deploy a server that receives HTTP requests from a client. The server hosts the developed API, which processes input prompts and returns the model’s output, specifically the generated decoded tokens. The experiment involves selecting one of the seven serving configurations and performing the following steps for each configuration: for each model, the model is loaded, and inference is conducted on all 164 requests using the input dataset. During the inference process for each serving configuration, we collect the dependent variables using the two profilers described in Section 4.6. To ensure stability, we run the server for five minutes before starting the inferences and for another five minutes after completing the inferences with each serving configuration. The experiment is repeated 10 times to ensure consistency in the results, as this number of repetitions was sufficient to observe minimal variability and provide reliable medians across configurations. Still, each repetition of the experiment involves 7 serving configurations, 12 models, and 164 input requests per model, resulting in a substantial amount of data collected.
@@ 4.8 Data analysis
RQs Variable Statistical test/
Post-hoc test Null hypothesis
@@ 5 Results
In this section, we report the results of our experiment, addressing our research questions and highlighting key takeaways. We initially evaluated seven deep learning serving configurations. Following the experiments, we reviewed the outputs of the scripts and inferences, identifying error-prone responses in the JIT configurations, $\langle JIT,CPU\rangle$ and $\langle JIT,CUDA\rangle$. JIT’s Trace mode generated error-prone responses because it records a fixed computational graph from sample inputs and cannot account for conditions outside the traced path. Given these issues, and to preserve fairness in our comparisons, we excluded JIT configurations ($\langle JIT,CPU\rangle$ and $\langle JIT,CUDA\rangle$) from further analysis. The results presented here focus on comparisons across the final five selected configurations while using results from twelve different code generation SLMs, as shown in Tables 45 and Figures 35 (one color for each model in the box plots).
In some figures, as in Figure 3, we grouped four different models in size category 3, each with its own mean values. That grouping inevitably produces greater overall dispersion than looking at any one of those models separately. Consequently, the difference stems from aggregating models with distinct baselines and performance profiles into a single category rather than reflecting instability in our measurements. As an example, Figures 68 present box plots for the model pythia-410m, shown before aggregation by model size category. These additional figures illustrate execution time, energy consumption, and computing-resource utilization across all serving configurations for a single model, showing the low variability observed during the experiment. Furthermore, the replication package contains disaggregated plots for all individual models, which consistently exhibit similarly low variability.
Due to the extensive amount of experimental data generated from assessing twelve distinct models across five serving configurations, we aggregated results into clearly defined model-size categories (see Table 1). The categories are organized based on model size category, with category 1 representing the smallest models and category 3 representing the largest models. Median values were specifically chosen to summarize categories due to their robustness against outliers, given substantial variability observed within size categories. Meanwhile, mean values were employed to summarize serving configurations owing to their more uniform distributions, facilitating direct comparative analysis across configurations.
It is important to explicitly note that the detailed model-by-model analysis was primarily conducted at the model level (without aggregation), and remains fully available in our replication package. Aggregated data here serves solely as a concise, high-level representation of overall trends.
5.1 Energy consumption of code generation SLMs(RQ1).
@@
The median energy consumption for configurations using the CUDA execution provider is consistently lower than that of the CPU execution provider. Notably, the $\langle TORCH,CUDA\rangle$ configuration stands out as the most energy-efficient across all models, a trend that is also evident in the aggregated results.
The $\langle ONNX,CUDA\rangle$ configuration exhibits the highest median energy consumption values among the CUDA execution provider configurations. Similarly, $\langle TORCH,CPU\rangle$ configurations consume the most energy across the majority of models when using the CPU execution provider. This pattern is also reflected in the aggregated results.
ONNX optimizations are significantly more effective when paired with the CPU execution provider, demonstrating better energy efficiency compared to the deep learning framework alone. Among the optimized runtime engines, ONNX along with OV, ONNX emerges as the most energy-efficient in CPU execution provider configurations. However, when using the CUDA execution provider, ONNX fails to provide the same level of energy savings, as it does not outperform the baseline TORCH runtime engine, which remains the most energy-efficient. This discrepancy in energy performance likely stems from the way the ONNX runtime engine interacts with each execution provider.
Although execution time alone is not a sufficient proxy for energy consumption, as discussed later, the substantial differences observed between CPU and CUDA configurations suggest that extended inference durations likely contribute to increased energy use in CPU-based settings (see Table 4). This prolonged inference time, especially with more complex models, may cause CPU configurations to use substantially more energy. The inefficiency of the CPU execution provider in energy conservation highlights the need for optimized resource allocation and management in deep learning serving environments to minimize unnecessary energy consumption. In contrast, CUDA configurations, despite utilizing both CPU and GPU resources, were far more energy-efficient for intensive tasks due to their significantly faster execution times.
@@
Answer to RQ1: Energy consumption is significantly affected by the serving configuration, with CUDA configurations being more efficient than CPU configurations. $\langle TORCH,CUDA\rangle$ provides the best energy efficiency, while optimized runtime engines like OV and ONNX, particularly ONNX, reduce energy consumption in CPU configurations but are less effective with CUDA compared to TORCH.
@@
1 $\langle ONNX,CPU\rangle$ 45069.35 655.09
1 $\langle OV,CPU\rangle$ 58374.88 561.61
1 $\langle TORCH,CPU\rangle$ 64993.75 773.27
2 $\langle ONNX,CPU\rangle$ 178641.05 2100.83
2 $\langle OV,CPU\rangle$ 223821.62 1355.18
2 $\langle TORCH,CPU\rangle$ 216854.71 2583.76
3 $\langle ONNX,CPU\rangle$ 393187.23 3958.18
3 $\langle OV,CPU\rangle$ 400527.50 2300.14
3 $\langle TORCH,CPU\rangle$ 512731.24 5668.91
1,2,3 $\langle ONNX,CPU\rangle$ 205632.54 2238.03
1,2,3 $\langle OV,CPU\rangle$ 227574.67 1405.64
1,2,3 $\langle TORCH,CPU\rangle$ 264859.90 3008.65
@@
1 $\langle ONNX,CUDA\rangle$ 13404.99 20692.81 34138.28 277.12
1 $\langle TORCH,CUDA\rangle$ 5193.32 11250.87 16432.53 124.03
2 $\langle ONNX,CUDA\rangle$ 32113.95 55652.21 87788.33 630.64
2 $\langle TORCH,CUDA\rangle$ 7287.23 24247.47 31525.72 168.74
3 $\langle ONNX,CUDA\rangle$ 39012.21 81516.82 120509.17 763.59
3 $\langle TORCH,CUDA\rangle$ 12196.07 50052.11 62301.62 260.62
1,2,3 $\langle ONNX,CUDA\rangle$ 28177.05 52620.61 80811.93 557.12
1,2,3 $\langle TORCH,CUDA\rangle$ 8225.54 28516.82 36753.29 184.46
5.2 Execution time of code generation SLMs (RQ2).
@@
Extreme cases are observed with models such as phi2 using $\langle TORCH,CPU\rangle$, where the median execution time (5983.52 seconds) is more than 20 times longer than with the $\langle TORCH,CUDA\rangle$ configuration (251.97 seconds), highlighting significant bottlenecks in CPU-based processing for complex computational tasks. Similar extreme cases are noted within the largest model size category (Category 3) using TORCH, where the CPU execution provider’s execution time (5668.91 seconds) exceeds the CUDA execution provider’s time (260.62 seconds) by more than 20 times.
@@
Our experiments revealed that while execution time optimization often aligns with reduced energy consumption, there exist notable exceptions. Specifically, within CPU configurations, the $\langle OV,CPU\rangle$ configuration exhibited faster execution times compared to $\langle ONNX,CPU\rangle$ but consistently required higher energy consumption. This explicitly shows scenarios where optimizing time performance alone may not directly lead to energy efficiency.
Welch’s ANOVA confirmed significant differences in execution times across different runtime engines. The subsequent Games-Howell post-hoc tests annotate that all configurations are statistically distinct from one another, confirming the marked impact of the execution provider and runtime engine choice on the efficiency of computational tasks (H.2.0). The only exceptions were that the execution times of the $\langle ONNX,CPU\rangle$ and $\langle TORCH,CPU\rangle$ configurations did not differ significantly for three out of the twelve models (bloomz-560m, codeparrot-small, and tiny-starcoder), resulting in the rejection of H.2.0 for only these models.
Answer to RQ2: Execution time is significantly influenced by the serving configuration, with CUDA configurations consistently outperforming CPU configurations. $\langle TORCH,CUDA\rangle$ delivers the best performance, while OV and ONNX optimize execution time, especially OV, with CPU execution provider but do not achieve comparable improvements with CUDA execution provider, reflecting similar patterns seen in energy consumption.
5.3 Computing-resource utilization of code generation SLMs (RQ3).
@@
For statistical evaluation, different tests were applied depending on the normality and variance homogeneity of the data. Welch’s ANOVA was used for data with unequal variances but normal distributions, while the Kruskal-Wallis test was employed for non-normally distributed data. The significance of differences between configurations was assessed using these tests, with p-values indicating significant differences with values below $\alpha=0.05$. Furthermore, post-hoc analyses using Games-Howell or Dunn’s test were performed to identify specific pairwise differences. In terms of computing-resource utilization metrics, all serving configurations showed significant differences from one another (H.3.1.0, H.3.2.0, H.3.3.0, H.3.4.0, and H.3.5.0), with a few exceptions. Specifically, RAM usage between certain serving configuration pairs—such as $\langle OV,CPU\rangle$ and $\langle TORCH,CPU\rangle$, $\langle ONNX,CUDA\rangle$ and $\langle TORCH,CUDA\rangle$, $\langle OV,CPU\rangle$ and $\langle ONNX,CPU\rangle$, as well as $\langle TORCH,CPU\rangle$ and $\langle ONNX,CUDA\rangle$—did not exhibit significant differences in two out of the twelve models for each pair. As a result, H.3.2.0 was rejected for those specific cases.